OpenStudy (anonymous):

find the nth term 1 2 3 4 5 6 2 8 18 32 50 ? show work!

OpenStudy (amistre64):

which nth term :)

OpenStudy (anonymous):

6

OpenStudy (anonymous):

1=2, 2=8, 3=18, 4=31 5=50 6=?

OpenStudy (amistre64):

2 = 2 2+6 = 8 8+10 = 18 18+14 = 32 not a pattern there

OpenStudy (amistre64):

got it

OpenStudy (chaise):

n = (n-1)+1; n1 = 1

OpenStudy (amistre64):

an + 2(1,3,5,7,9,...)

OpenStudy (anonymous):

Terms : 2 8 18 32 50 72 Diffs : 6 10 14 18 22 2nd diffs : 4 4 4 4

OpenStudy (amistre64):

well, if we start at: a1 = 0 + 2(1) = 2 a2 = 2 + 2(3) = 8 a3 = 8 + 2(5) = 18 a4 = 18 + 2(7) = 32

OpenStudy (anonymous):

so whats the nth term?????

OpenStudy (amistre64):

a5 = 32 + 2(9) 32 + 18 50 woohoo!!!

OpenStudy (anonymous):

do you want the 6th or the nth ?

OpenStudy (anonymous):

nth

OpenStudy (anonymous):

the 6th is 72

myininaya (myininaya):

joe is always really good at these

OpenStudy (amistre64):

an = 2(2n-1) maybe :)

OpenStudy (anonymous):

These problems dont make sense to me anymore =/

myininaya (myininaya):

lol

OpenStudy (anonymous):

but whats the nth!

OpenStudy (anonymous):

There is an algorithm for this . Just answered a similar Q about it - hold on.

OpenStudy (amistre64):

an + 2(2n-1) since the odds are 2k -1

OpenStudy (chaise):

We can use the formula tn = a+(n-1)d Where a = first number in the equation n = term number d = common difference Solve for n

OpenStudy (amistre64):

almost got my brain around it lol

OpenStudy (anonymous):

to follow what the person who make the problem wants us to do, since the 2nd differences are constantly 4, its a quadratic polynomial in the form an^2+bn+c

OpenStudy (amistre64):

oh whats the fun in using premade wheels ... i like to reinvent them as i go ;)

OpenStudy (anonymous):

you can plug in values of n (1, 2, 3,...) and create a system of equations.

myininaya (myininaya):

n(2n) or 2n^2

myininaya (myininaya):

so we see 2=2(1) 8=2(2(2)) 18=3(2(3)) 32=4(2(4)) so on....

OpenStudy (amistre64):

an = an-1 + 2(2n-1) = an-2 + 2(2n-1) + 2(2n-1) = an-3 + 2(2n-1) + 2(2n-1)+ 2(2n-1) myin seems to be onto something easier lol

OpenStudy (anonymous):

for the record :) http://www.mathsisfun.com/algebra/sequences-finding-rule.html

OpenStudy (anonymous):

so the nth term is an-3+2(2n-1)+2(2n-1)+2(2n-1)???? and thanks! haha

OpenStudy (amistre64):

i wish there was set "rules" :)

OpenStudy (anonymous):

its just 2n^2.

OpenStudy (amistre64):

it prolly simplifies out to myins

OpenStudy (amistre64):

a1 = 2.1.1 a2 = 2.2.2 a3 = 2.3.3 an = 2.n.n = 2n^2

OpenStudy (anonymous):

divide those terms by 2, they are perfect squares. thats probably what the creator of the problem intended. There are infinitely many formulas that could work though, and infinitely many "6th terms"

OpenStudy (anonymous):

And Joe is right :) http://www.wolframalpha.com/input/?i=2+++++8++++++18++++++32+++++50

myininaya (myininaya):

i seen n(2n) and then the simplification or product is 2n^2 way to multiply joe! thats why you are one of my favorites ;)

OpenStudy (amistre64):

what would you do-o-o 4 a joemath bar lol

OpenStudy (anonymous):

lol <.<

OpenStudy (anonymous):

this link will work: http://www.wolframalpha.com/input/?i=2+8++18++32++50

myininaya (myininaya):

i want a heath bar with some vanilla icecream

OpenStudy (amistre64):

the links are on the fritz by the way

OpenStudy (anonymous):

give me 5 numbers between -5 and 5, without picking 0.

OpenStudy (amistre64):

.031172 is my first pick

OpenStudy (anonymous):

hahah thanks everyone!

OpenStudy (anonymous):

i guess i should specify integers, to make the calculations easier on myself lol

OpenStudy (amistre64):

;)

OpenStudy (anonymous):

1 2 3 4 5

OpenStudy (amistre64):

3 1 1 4 2 are 5 nukbers

OpenStudy (amistre64):

...the k key is my new nemsis

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

alright, give me a couple of minutes, i'll pop out another formula that fits those first 5 numbers for n = 1, 2, 3, 4, 5, and gives some other 6th term.

OpenStudy (amistre64):

"between -5 and 5 tends to mean up to but not including ... mathically that is

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-14%2F3%281%29^n%2B10%2F3%282%29^n%2B1%2F3%283%29^n-1%2F3%284%29^n%2B1%2F15%285%29^n%2C+n+%3D+1 you can change the value of n at the end of the equation to verify. According to this formula, the 6th term is 128 lol

OpenStudy (anonymous):

im actually thinking about writing an undergraduate thesis on this idea. <.<